prxncessjamyria:

Initially, there were only 197 weeds at a park. The weeds grew at a rate of 25% each week. The following function represents the weekly weed growth: f(x) = 197(1.25)x. Rewrite the function to show how quickly the weeds grow each day and calculate this rate as a percentage.

Babyplier:

@mercury

Babyplier:

@dude

Babyplier:

@hero

Mercury:

next time please post the answer choices A.) f(x) = 197(1.25)^7x; grows at a rate of approximately 2.5% daily B.) f(x) = 197(1.25^7)^x; grows at a rate of approximately 4.77% daily C.) f(x) = 197(1.03)^x; grows at a rate of approximately 0.3% daily D.) f(x) = 197(1.03)^7x; grows at a rate of approximately 3% daily since it grows 25% per week (7 days), it grows 25%/7 per day. change the growth rate to reflect this change. now, since we want to re-write the equation in terms of days, we also need to multiply the exponent (which is currently in weeks) by 7. perform these changes on the original equation to get the new equation.